OpenStudy (anonymous):

Find the general solution of the linear equation: y'=1+x+y+xy Not sure what the P(x) is when written as y'+P(x)=Q(x)

OpenStudy (anonymous):

WHY THE FREAK DID U POST IT A THOUSAND TIMES!?!?!?

OpenStudy (anonymous):

this is my first time posting this question

OpenStudy (dumbcow):

rewrite \[\frac{dy}{dx}=y(x+1) +(x+1) = (x+1)(y+1)\] separate variables \[\frac{dy}{y+1} = (x+1) dx\] integrate both sides , then solve for y

OpenStudy (dumbcow):

you don't need to use y' +P(x)y = Q(x) since this is separable

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks